Find the length of a carpet 75 cm wide to cover the floor of a room whose dimensions are 8 meters by 6 meters.
Answer by bucky(2189)

The floor of the room is 8 meters by 6 meters. The area of this floor is the length multiplied
by the width. In equation for this is:
.
A = L*W
.
So the area is:
.
A = 8 * 6 = 48 square meters.
.
We are told that the carpet width is 75 cm. Be careful. We have to change that to meters.
The carpet is 0.75 meters wide.
.
The length of the carpet times its width must equal 48 square meters. Return to the area
equation:
.
A = L*W
.
We know the area we want is 48 square meters. We also know that the carpet width is 0.75 meters.
So we substitute those two values into the Area equation to get:
.
48 = L * 0.75
.
Divide both sides of this equation by 0.75 meters to get:
.
48/0.75 = L
.
If you do the division you will find that the length of carpet needed is 64 yards.
.
Hope this helps you to understand the problem.